The yield to maturity on 1-year zero-coupon bonds is currently 7%; the YTM on 2-year zeros is 8%.

The Government plans to issue a 2-year maturity coupon bond, paying coupons once per year with a

coupon rate of 9%. The face value of the bond is $100.

a. At what price will the bond sell?

b. What will the yield to maturity on the bond be? (Hint: Use a financial calculator to get the YTM)

c. If the expectations theory of the yield curve is correct, what is the market expectation of the price that

the bond will sell for next year?

d. Recalculate your answer to (c) if you believe in the liquidity preference theory and you believe that the

liquidity premium is 1%.
Answer #1

a) Bond Price can be calculated using PV function on a calculator

N = 2, PMT = 9% x 100 = 9, FV = 100, I/Y = 8%

=> Compute PV = $101.78

b) YTM on the bond is the same as that on 2 year zero coupon = 8% can also be calculated using the I/Y function on a calculator

c) Using expectation theory,

(1 + s2)^2 = (1 + s1) x (1 + 1f1)

=> (1 + 8%)^2 = (1 + 7%) x (1 + 1f1)

=> 1f1 = 9.01% is the one year rate, one year later.

Price of the bond a year later = (PMT + FV) / (1 + 1f1) = (9 + 100) / 1.0901 = $100